the geometric length of a bar magnet is 24cm. the length of the magnet is
 

Grade:12

6 Answers

Nikhil Upadhyay
8 years ago
the magnetic length of a magnet is aproximately 6\7 of its geometric length
therefore the answer is 20.57cm(approx) 
Anoopam Mishra
8 years ago
Magnetic length = 5/6 of Geometric length = 5/6 * 24 = 20 cm
Anoopam Mishra
8 years ago
Magnetic length is about 5/6 times or 83.3 % of the geometric length and not 6/7 times as answered by Nikhil.
Nikhil Upadhyay
8 years ago
Magnetic Length is 0.85 times the geometric length thus it is 6/7 times of the geometric length
Anoopam Mishra
8 years ago
Nikhil search on Google, most of them say that it is 5/6 only few say that it is 6/7.
Shamanth
7 years ago
Magnetic length is 5/6 of geometric length if you calculate magnetic length for 24 cm of geometric length you will get the answer as 20 cm..I hope you liked it .